OpenStudy (waheguru):

How is this a function y= √(x+2 the textbook says its a function but I thought when we square root a number the answers can be either a positive or a negative

OpenStudy (anonymous):

use vertical line test http://www.wolframalpha.com/input/?i=plot+sqrt%28x%2B2%29

OpenStudy (waheguru):

Yea your right ! thanks =)

OpenStudy (waheguru):

but when you plug in an x value we get 2 y values ?

OpenStudy (waheguru):

Say we use x = 2 then then y could possibly be +2 or -2?

OpenStudy (waheguru):

@cinar

OpenStudy (anonymous):

no not like that..

OpenStudy (waheguru):

but when we square root something we can either have a postiive or a negative

OpenStudy (anonymous):

if x^2=4, then x=-+2 but here if x=2, then y=2

OpenStudy (waheguru):

I get the first part but...

OpenStudy (waheguru):

@nincompoop

OpenStudy (waheguru):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

cinar is absolutely correct there are two solutions to x^2 = 4, and they are x = 2 or x = -2 but the square root function is a function because plugging in any nonnegative real number for x produces EXACTLY ONE nonnegative real number for y

jimthompson5910 (jim_thompson5910):

so yes x^2 = k has two solutions x = plus/minus sqrt(k) but the square root function itself produces one output per input, and this output is either 0 or positive if you want negative outputs, you have to force it to be negative by sticking a negative out front like this: -sqrt(k)